Let R be a ring, I an ideal of R, J an ideal of I. Show that if J has a unity, then J is

an ideal of R. Also give an example to show that if J does not satisfy this condition

it need not be an ideal of R.


1
Expert's answer
2021-06-18T04:43:58-0400

A)

Let "j \in J" and "r\in R".

Also, let e be the unity of J.

Then, "re \in I" (since I is an ideal of R and e is also in I)

And so "(re)j \in J" (since J is an ideal of I)

But "(re)j = rj"

i.e "rj \in J"

Hence, J is an ideal of R.

B) Consider "\mathbb{Z_4}" and "\{0,2\}"

Let R = "\mathbb{Z_4}" , I "= \mathbb{Z_4}" and J = "\{0,2\}"

We see from this example that the condition above will not be satisfied because, J has defined here is only a sub ring and not an ideal.